OpenStudy (anonymous):

What are the solutions of x2 + 3x − 18 = 0? Show all work for full credit

OpenStudy (mathstudent55):

Do you know how to factor?

OpenStudy (anonymous):

sadly no

OpenStudy (mathstudent55):

You need to factor the left side which is a three term polynomial, also called a trinomial.

OpenStudy (anonymous):

I know i shouldnt do this but i was wondering if you could just give me the answer its my last question and i have to run to the store

OpenStudy (mathstudent55):

A trinomial of the form you have: x^2 + ax + b if it's factorable will factor into two binomials of the form (x + p)(x + q) where p times q = b and p + q =a

OpenStudy (anonymous):

so my parents are rushing me

OpenStudy (mathstudent55):

You need two numbers that multiply to -18 and add to 3. They are 6 and -3 (x + 6)(x -3) = 0 x + 6 = 0 or x - 3 = 0 x = -6 or x = 3

OpenStudy (anonymous):

ok thank you so much